OpenStudy (anonymous):

find solution in the interval [0,2pi): sin(2x)= sqrt3/2

OpenStudy (anonymous):

do you have a link?

OpenStudy (bluemiku):

there's no link?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i can rewrite it for you:\[find solution \in the interbal [0,2\pi): \sin(2x)=\frac{ \sqrt{3}}{ 2 }\]

OpenStudy (anonymous):

whoops, it supposed to say find in the intervals [0,2pi). Not sure what happened when I wrote the equation lol

jimthompson5910 (jim_thompson5910):

If you used the unit circle, what is the value of t that makes sin(t) = sqrt(3)/2 true?

OpenStudy (anonymous):

11pi/6

jimthompson5910 (jim_thompson5910):

11pi/6 is in Q4 where sine is negative

OpenStudy (anonymous):

pi/6?

jimthompson5910 (jim_thompson5910):

closer but no sin(pi/6) = 1/2

jimthompson5910 (jim_thompson5910):

we want the sine value to be sqrt(3)/2

OpenStudy (anonymous):

pi/3

jimthompson5910 (jim_thompson5910):

good, sin(pi/3) = sqrt(3)/2

jimthompson5910 (jim_thompson5910):

what other value of t makes sin(t) = sqrt(3)/2 true?

OpenStudy (anonymous):

2pi/3

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so if sin(2x) = sqrt(3)/2 then 2x = pi/3 or 2x = 2pi/3

jimthompson5910 (jim_thompson5910):

add on "+2pi*n" to capture all coterminal angles if sin(2x) = sqrt(3)/2 then 2x = pi/3+2pi*n or 2x = 2pi/3+2pi*n

jimthompson5910 (jim_thompson5910):

n is any integer

OpenStudy (anonymous):

ok, thank you!

jimthompson5910 (jim_thompson5910):

what is the next step

OpenStudy (anonymous):

oh theres more? lol..hmm, I don't know. This is the first time I've had a 2x in sine.

OpenStudy (anonymous):

divide by 2?

jimthompson5910 (jim_thompson5910):

you have 2x = .... you want x = ...

jimthompson5910 (jim_thompson5910):

yes you divide everything by 2

jimthompson5910 (jim_thompson5910):

or better yet, multiply everything by 1/2

jimthompson5910 (jim_thompson5910):

2x = pi/3+2pi*n (1/2)*2x = (1/2)*(pi/3)+(1/2)*2pi*n x = pi/6 + pi*n

OpenStudy (anonymous):

pi/6+pi*n and 2pi/6+pi*n?

jimthompson5910 (jim_thompson5910):

2x = 2pi/3+2pi*n (1/2)*2x = (1/2)*(2pi/3)+(1/2)*2pi*n x = pi/3 + pi*n

OpenStudy (anonymous):

ohh! Ok.

jimthompson5910 (jim_thompson5910):

if sin(2x) = sqrt(3)/2 then 2x = pi/3+2pi*n or 2x = 2pi/3+2pi*n that turns into x = pi/6 + pi*n or x = pi/3 + pi*n where n is any integer

jimthompson5910 (jim_thompson5910):

you need to now generate particular solutions by plugging in values of n make sure that x is in the interval [0, 2pi)